OpenStudy (anonymous):

Complete the square to solve the equation. x^2+8x-5=15

OpenStudy (anonymous):

x^2 + 8x - 20 =0 x^2 + 2*4*x - 20 + 16 -16=0 (x + 4)^2= 36 x + 4 = (36)^1/2 x + 4 = |6| x = -10,2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

some weird business here. answer is right though

OpenStudy (anonymous):

try \[x^2+8x=20\] \[(x+4)^2=20+16=36\] now next line should be \[x+4=6\] or \[x+4=-6\] not \[x+4=|6|\]

OpenStudy (anonymous):

as i said, answer is correct but method is not right

OpenStudy (anonymous):

Thank you satellite and ishan!